### v1

# Operational meaning of the fidelity and the Fisher-Bures path length

Abstract
We show that the minimum entropy production in near-reversible quantum state transport along a path is simple function of the path length measured according to the Fisher–Bures metrics. Consequently, the minimum entropy production along a geodesics is the simple function of fidelity. Hence the sharp values of path lengths, also called statistical lengths, and of fidelities obtain operational meaning to quantify the residual irreversibility in near-reversible state transport. Compared to fidelity’s operative interpretations based on Uhlmann’s theorem, our proposal does not assume extended operations on the purified states. As to the classical limit, the Bhattacharryya fidelity obtains a sharp operational meaning after eighty years.
I Introduction
A quantitative comparison between two probability distributions $p_{\alpha}$ and $q_{\alpha}$ can be based on the Bhattacharyya coefficient published in 1945 Bhattacharyya (1943):
$F(p,q)=\sum_{\alpha}\sqrt{p_{\alpha}q_{\alpha}}~{}\in[0,1].$ (1)
This function defines a measure $F\in[0,1]$ of closeness between $p$ and $q$. Fifty years later, to measure the closeness of two quantum states $\hat{\rho}$ and $\hat{\sigma}$, Jozsa borrowed the expression known in mathematics Bures (1969); Uhlmann (1976) and called it fidelity Jozsa (1994):
$F(\hat{\rho},\hat{\sigma})=\|\sqrt{\hat{\rho}}\sqrt{\hat{\sigma}}\|_{\mathsf{tr}}=\mathsf{tr}\sqrt{\sqrt{\hat{\sigma}}\hat{\rho}\sqrt{\hat{\sigma}}}~{}\in[0,1].$ (2)
Quantum fidelity is the central tool in modern quantum informatics Nielsen and Chuang (2001). In the special (classical) case when $[\hat{\rho},\hat{\sigma}]=0$, we can write $\hat{\rho}=\mathsf{diag}[p_{1},p_{2},\dots,p_{\alpha},\dots]$ and $\hat{\sigma}=\mathsf{diag}[q_{1},q_{2},\dots,q_{\alpha},\dots]$. The quantum fidelity (2) coincides with the classical fidelity (1).
One of our goals is the operational interpretation of $F(q,p)$ because it is not known Nielsen and Chuang (2001). For the quantum fidelity $F(\hat{\rho},\hat{\sigma})$ a simple operational interpretation was already proposed Dodd and Nielsen (2002), based on Uhlmann’s theorem Uhlmann (1976), and goes like this. Assume that $\hat{\rho}$ and $\hat{\sigma}$ are reduced states of two pure states, respectively, in a larger system. Assume that we have access to this larger system. Then the distinguishability of $\hat{\rho}$ and $\hat{\sigma}$ is at least as high as the distinguishability of two arbitrary pure states whose overlap (modulus of scalar product) is given by the fidelity $F(\hat{\rho},\hat{\sigma})$. This interpretation assumes the operation on the larger system where both $\hat{\rho}$ and $\hat{\sigma}$ are pure. Various theorems contain the sharp value of fidelity but their meaning should always include operations on the ‘environmental’ quantum system.
Fidelities are related to distances, also called Fisher–Bures statistical distances, on the underlying Riemannian geometries. The infinitesimal distance $d\ell$ defines the Fisher metrics Fisher (1925) between the probability distributions $p$ and $p+dp$ and the Bures metrics Bures (1969) between the quantum states $d\hat{\rho}$ and $\hat{\rho}+d\hat{\rho}$:
$(d\ell)_{Fisher}^{2}$ $=$ $2\sum_{\alpha}(d\sqrt{p_{\alpha}})^{2}=\sum_{\alpha}\frac{(dp_{\alpha})^{2}}{p_{\alpha}},$ (3)
$(d\ell)_{Bures}^{2}$ $=$ $2\mathsf{tr}(d\sqrt{\hat{\rho}})^{2}=\mathsf{tr}\left(d\hat{\rho}\frac{2}{\hat{\rho}_{L}+\hat{\rho}_{R}}d\hat{\rho}\right).$ (4)
The labels $L/R$ mean $\hat{\rho}_{L}$ acts from the left, as usual, while $\hat{\rho}_{R}$ acts from the right. The local Fisher metrics is the special case of the local Bures metrics when $[\hat{\rho},d\hat{\rho}]=0$. Consider the geodesics between $p$ and $q$ or between $\hat{\rho}$ and $\hat{\sigma}$. The geodesic distances $\ell$ are functions of the respective fidelities:
$\ell(p.q)_{Fisher}$ $=$ $2\arccos F(p,q)~{}\in[0,\pi],$ (5)
$\ell(\hat{\rho}.\hat{\sigma})_{Bures}$ $=$ $2\sqrt{1-F^{2}(\hat{\rho},\hat{\sigma})}~{}\in[0,2].$ (6)
As we said, when $\hat{\rho}$ and $\hat{\sigma}$ commute, then $F(q,p)$ is the special case of $F(\hat{\rho},\hat{\sigma})$ and the local Fisher geometry (3) is the special case of the local Bures geometry (4). This is no longer the case with the above global distances since $\ell_{Bures}=2\sin(\tfrac{1}{2}\ell_{Fisher})$. The Bures geodesic distance is shorter than the Fisher geodesic distance: The latter is the minimum length of paths but through the commuting states.
The infinitesimal distances, i.e.: the local metrics, have exact meaning in metrology via the famous Cramér-Rao Cramér (1946); Radhakrishna Rao (1945) and quantum Cramér-Rao theorems Braunstein and Caves (1994). Suppose an unknown state —or probability distributions $p$ in the classical case— at the unknown length $\ell$ measured from one end of a known path. We perform $N$ independent measurements on the unknown state to estimate $\ell$, i.e., to estimate the unknown state. If $N\rightarrow\infty$, optimum measurements ensure that the mean squared error of estimation goes to zero as $(\Delta\ell)^{2}=1/N$. This interpretation concerns the local geometry at one point and does not tell us anything about non-local features like generic Fisher–Bures distances.
The present work proposes the operational meaning of path lengths in the Fisher–Bures metrics, which implies the meaning of the geodesics and the fidelities. Differently from ref. Dodd and Nielsen (2002), we are not relying on Uhlmann’s theorem and the ‘environmental’ system. The concept is that in near-reversible transport along a path $\gamma$ between two predefined states $\hat{\rho},\hat{\sigma}$ the minimum entropy production is quantified by the length $\ell_{\gamma}(\hat{\rho},\hat{\sigma})$.
II State transport by equilibrating reservoirs
In preparation, we define the transport of the system’s initial state $\hat{\rho}$ into the final state $\hat{\sigma}$ in contact with a single reservoir. We use the reservoir model of ref. Diósi et al. (2006) and the theorem therein (cf. Csiszár et al. (2007) for the rigorous proof).
The reservoir consists initially of $n$ independent systems in state $\hat{\sigma}$ each and we start from the composite system-reservoir state $\hat{\rho}\otimes\hat{\sigma}^{\otimes n}$. The contact of the system and the reservoir assumes a reversible step, a swap between the system’s state $\hat{\rho}$ and (e.g. the 1st) one of the reservoir’s states $\hat{\sigma}$:
$\hat{\rho}\otimes\hat{\sigma}^{\otimes n}\Rightarrow\hat{\sigma}\otimes\hat{\rho}\otimes\hat{\sigma}^{\otimes(n-1)}.$ (7)
The second step, concerning the reservoir only, is the irreversible relaxation of the reservoir to a homogeneous state, modeled by the following twirl over the permutation group:
$\hat{\rho}\otimes\hat{\sigma}^{\otimes(n-1)}\Rightarrow\frac{1}{n}\sum_{k=0}^{n-1}\hat{\sigma}^{\otimes k}\otimes\hat{\rho}\otimes\hat{\sigma}^{\otimes(n-k-1)}.$ (8)
According to this model, the transport $\hat{\rho}\rightarrow\hat{\sigma}$ of the system state leads to the irreversible relaxation inside the reservoir. In the infinite reservoir limit $n\rightarrow\infty$, fortunately, the entropy production of relaxation can be expressed as the relative entropy $S(\hat{\rho}\|\hat{\sigma})=\mathsf{tr}(\hat{\rho}\ln\hat{\rho}-\hat{\rho}\ln\hat{\sigma})$:
$\!\!\!\!\!S\!\left(\frac{1}{n}\sum_{k=0}^{n-1}\hat{\sigma}^{\otimes k}\otimes\hat{\rho}\otimes\hat{\sigma}^{\otimes(n-k-1)}\right)-S\!\left(\hat{\rho}\otimes\hat{\sigma}^{\otimes(n-1)})\right)$
$n\rightarrow\infty]{}S(\hat{\rho}\|\hat{\sigma}).$ (9)
This is the entropy production of the single-step state transport as well:
$\Delta S=S(\hat{\rho}\|\hat{\sigma}).$ (10)
The entropy production can be smaller and can even go to zero if we apply sequential equilibration with intermediate reservoirs along a path, such that the stepsizes go to zero. The following feature of relative entropy will be important:
$S(\hat{\rho}|\hat{\rho}+d\hat{\rho})=\frac{1}{2}(d\ell)^{2},$ (11)
where $d\ell$ is defined by the Bures metrics (4). Consider a smooth (not necessarily geodesic) path $\gamma$ from $\hat{\rho}$ to $\hat{\sigma}$. We use $N$ reservoirs to transport $\hat{\rho}$ into $\hat{\sigma}$ in $N$ steps. Consider a monotone sequence of intermediate states $\hat{\rho}_{i}$ along the path, such that $\hat{\rho}=\hat{\rho}_{0}$ and $\hat{\rho}_{N}=\hat{\sigma}$. Perform the transports $\hat{\rho}\Rightarrow\hat{\rho}_{1}\Rightarrow\hat{\rho}_{2}\Rightarrow\dots\Rightarrow\hat{\rho}_{N-1}\Rightarrow\hat{\sigma}$. The total entropy production is the sum of the $N$ yields:
$\Delta S=\sum_{i=0}^{N-1}S(\hat{\rho}_{i}\|\hat{\rho}_{i+1}).$ (12)
In the limit $N\rightarrow\infty$ each step $\Delta\ell_{i}$ between $\hat{\rho}_{i}$ and $\hat{\rho}_{i+1}$ can go to zero. In this asymptotic regime, applying eq. (11) we can thus write
$\Delta S=\frac{1}{2}\sum_{i=0}^{N-1}(\Delta\ell_{i})^{2}.$ (13)
This is an important expression but the r.h.s. is not yet the function of the total path length $\ell_{\gamma}(\hat{\rho},\hat{\sigma})$ while the lengths elements are constrained by it:
$\sum_{i=0}^{N-1}\Delta\ell_{i}=\ell_{\gamma}(\hat{\rho},\hat{\sigma}).$ (14)
Keeping this constraint, let us minimize the entropy production (13). It is trivial that each term on the r.h.s. must have the same value, i.e., $\Delta\ell_{i}=\ell_{\gamma}(\hat{\rho},\hat{\sigma})/N$ for all $i$. This means that the dense sequence $\hat{\rho}_{i}$ must be evenly distributed along the path, and the entropy production will be evenly distributed over the steps. The minimum of the total entropy production along the given path $\gamma$ takes this form:
$\Delta S_{\gamma}=\frac{\ell_{\gamma}^{2}(\hat{\rho},\hat{\sigma})}{2N}.$ (15)
The fastest approach to reversible transport is achieved on geodesics. Inserting the geodesic lengths (5-6) in (15) and its classical counterpart, the ultimate lower bounds on entropy production read
$\Delta S=\frac{2}{N}\left\{\begin{array}[]{ll}1-F^{2}&\mbox{(quantum)}\\ (\arccos F)^{2}&\mbox{(classical)}\end{array}\right.$ (16)
These express the desired sharp operational meaning of both the classical and quantum fidelities (1-2). The bounds are attainable in the limit $N\rightarrow\infty$ by operations on the system plus the reservoirs.
We find the simple meaning of the Bures length if we introduce the number $\nu$ of equilibrations per unit length, i.e., the density $\nu=N/\ell_{\gamma}$. Then the entropy production becomes the following linear function of the path length:
$\Delta S_{\gamma}(\ell)=\frac{\ell_{\gamma}}{2\nu}.$ (17)
This suggests the following operational meaning of the Bures length. Consider the near-reversible state transport via sequential equilibrations along a smooth path. If $\nu\gg 1$ is the density of equilibrations per unit path length then the entropy production is $(1/2\nu)$ per unit length.
III Remarks and summary
Paths in Fisher–Bures geometries are perhaps the simplest non-local objects and, as such, they can be interpreted as state transport processes. It has therefore seemed logical to look here for the meaning of the geometries. One should mention the standard Kantorovich–Wasserstein theory Kantorovich (1960); Vaserstein (1969) where the transport from distribution $p$ to $q$ happens via direct relocation of populations from the $p_{\alpha}$’s into the $q_{\beta}$’s. The transport of quantum state $\hat{\rho}$ into $\hat{\sigma}$, too, can work similarly after a single unitary rotation of $\hat{\rho}$ that makes it commute with $\hat{\sigma}$. This may seem much simpler than sequential equilibration. But the simple method assumes the detailed control of the states in question. If it is not possible because, e.g., the states are complicated many-body states then the method of equilibration becomes more valuable.
Thermodynamic systems are the typical examples. The Weinhold–Ruppeiner geometry Weinhold (1975); Ruppeiner (1979) defines the thermodynamic distances between thermodynamic states and the thermodynamic path lengths in general. It has long been known that the thermodynamic lengths coincide with statistical lengths Diósi et al. (1984) and they are rooted in the the Fisher geometry of Gibbs states Janyszek and Mrugała (1989). In thermodynamics, the expression $(\Delta\ell)^{2}/V$ quantifies how the statistical fluctuations of the thermodynamic parameters tend to zero in the thermodynamic limit of infinite volume $V\rightarrow\infty$. This does not interpret the global thermodynamic lengths $\ell$ but the local metrics. Studies of thermodynamic state transport Salamon and Nulton (1998); Salamon et al. (2023); Diósi (2002); Diosi et al. (1996) targeted a relationship between the thermodynamic path length and the thermodynamic entropy production. Salamon and Nulton Salamon and Nulton (1998) recognized that in discrete sequential equilibration the equilibration rates cancel from the expression of entropy production and they found the thermodynamic predecessor of eq. (15). Ref. Diósi (2002) was the first attempt at constructing an underlying classical microscopic mechanism of a specific irreversible state transport. The present work constructed the abstract microscopic generalization valid in the Fisher–Bures geometries of arbitrary quantum states including the classical special case of probability distributions. For this construction, the analytically tractable microscopic reservoir model and the related theorem, both proposed in ref. Diósi et al. (2006), have been instrumental.
The sharp, not just qualitative, operational (physical) meaning of classical and quantum fidelities has been demonstrated in the near-reversible transport of states. On the one hand, classical fidelity and the global Fisher statistical distance got an operational interpretation for the first time. On the other hand, quantum fidelity and Bures distance got an operational interpretation that works without operations on Uhlmann’s extension beyond the system of interest. The concrete choice of the reservoir and the equilibration protocol is not likely to be critical and alternative microscopic models should lead to the same relationships between entropy production and the Fisher–Bures geometry. Future investigations may replace our stepwise transport with some continuous protocol.
Acknowledgements.
This research was funded by the National Research, Development and Innovation Office (Hungary) Frontline Research Excellence Program (Grant No. KKP133827).